### v1

# A Statistical Learning Approach for Feature-Aware Task-to-Core Allocation in Heterogeneous Platforms

Abstract
Optimizing task-to-core allocation can substantially reduce power consumption in multi-core platforms without degrading user experience. However, many existing approaches overlook critical factors such as parallelism, compute intensity, and heterogeneous core types. In this paper, we introduce a statistical learning approach for feature selection that identifies the most influential features—such as core type, speed, temperature, and application-level parallelism or memory intensity—for accurate environment modeling and efficient energy optimization. Our experiments, conducted with state-of-the-art Linux governors and thermal modeling techniques, show that correlation-aware task-to-core allocation lowers energy consumption by up to 10% and reduces core temperature by up to 5C compared to random core selection. Furthermore, our compressed, bootstrapped regression model improves thermal prediction accuracy by 6% while cutting model parameters by 16%, yielding an overall mean square error reduction of 61.6% relative to existing approaches. We provided results based on superscalar Intel Core i7 12th Gen processors with 14 cores, but validated our method across a diverse set of hardware platforms and effectively balanced performance, power, and thermal demands through statistical feature evaluation.
Keywords Energy Optimization, Heterogeneous Platforms, Statistical Feature Evaluation, Task-to-Core Allocation
1 Introduction
Dynamic voltage and frequency scaling (DVFS) and task-to-core allocation are critical techniques for optimizing the performance and energy efficiency of embedded systems. Since power consumption is exponentially related to voltage and frequency, an effective DVFS strategy can reduce power consumption by up to 75% without impacting user experience Ratković et al. (2015). However, different workloads exhibit varying thermal and power consumption behaviors depending on task-to-core allocation strategies. In multiprocessor systems, cores with diverse characteristics, such as high-performance cores, low-power cores, and GPUs, offer unique performance trade-offs. Suboptimal core allocation can lead to thermal throttling, reduced reliability, and shorter device lifespan Hosseinimotlagh and Kim (2019). Furthermore, the cooling costs associated with overheating can account for approximately \$3 per watt of heat dissipation et al. (2003).
To address these challenges, existing approaches rely primarily on historical workload performance and sensor data to guide future processor behavior Maity et al. (2022). However, these approaches often fail to generalize in varying workload conditions and hardware architectures.
Multi-Core heterogeneous processors, which are widely used in automotive, telecommunications, and consumer electronics, must operate within strict power and thermal constraints while maintaining high performance. In heterogeneous architectures, DVFS and task-to-core allocation are essential for managing these constraints by optimizing the trade-off between power efficiency and performance. Feature selection plays a crucial role in identifying key system parameters, such as frequency, temperature, and voltage, which influence thermal stability and energy efficiency Shekarisaz et al. (2021). For example, in real-time automotive systems, DVFS strategies can minimize energy consumption without sacrificing response times, whereas in mobile devices, efficient task-to-core allocation extends battery life by utilizing low-power cores for background tasks.
Despite these advancements, the reliance on real-world data collection for policy optimization introduces computational overhead and inaccuracies due to hardware sampling delays. A promising alternative is the use of statistical learning models to infer system behavior from collected data, reducing reliance on extensive real-world testing. Properly trained models can capture the stochastic nature of embedded environments and provide accurate control signal generation Liu et al. (2021). This work proposes a statistical learning approach to develop efficient models using data augmentation and feature selection, thus enhancing energy-efficient task scheduling.
To the best of our knowledge, limited work has been done in applying statistical feature selection for DVFS policy optimization. Previous efforts, such as the use of extreme value theorem (EVT) to estimate upper bounds of energy consumption and execution time Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020), lack a comprehensive analysis of feature importance. Approaches such as XGBoost and decision trees have been employed to evaluate application-specific latency and energy features Liu et al. (2021); Sasaki et al. (2007), but a more robust, lightweight model suitable for multi-core platforms is needed.
This paper contributes the following:
  1. 1.
    A novel evaluation of feature importance for environment modeling and compressed learning for task-to-core allocation.
  2. 2.
    A correlation-aware task-to-core allocation approach to reduce temperature and energy consumption by strategically assigning tasks to uncorrelated cores.
  3. 3.
    A data augmentation strategy using bootstrapping to enhance model accuracy while reducing sample collection overhead.
  4. 4.
    Empirical validation against state-of-the-art thermal modeling methods, demonstrating up to 10% energy savings and $5^{\circ}$ temperature reduction. Our compressed regression model reduces prediction error by 6% while cutting the number of parameters by 16%, achieving a 61.6% reduction in mean square error.
2 Motivation and Challenges
Feature selection is a crucial data processing strategy in statistical learning, addressing challenges such as the curse of dimensionality, large data management, and performance unpredictability in multi-core platforms.
2.1 Curse of Dimensionality
Feature selection plays a vital role in mitigating the curse of dimensionality. As the number of features increases, the data space grows exponentially, leading to sparsity in high-dimensional spaces Li et al. (2017). This phenomenon can result in model overfitting, which degrades performance when predicting unseen data. In the context of heterogeneous systems, the increasing number and variety of processor cores—comprising performance, low-energy, and GPU cores, each characterized by frequency, temperature, and performance features—contribute to the expansion of the state space. As a result, efficient feature selection is necessary to reduce complexity and improve generalization.
2.2 Large Data Management
Feature selection is essential for handling the growing volume of data in DVFS and task-to-core allocation strategies. These strategies may rely on either static or streaming data, where static data consists of fixed historical samples, while streaming data continuously adapts to new input. With the increasing prevalence of streaming data, memory management becomes critical, as the unpredictable volume of incoming data can overwhelm system resources. Retaining unnecessary features significantly increases data storage and processing overhead, making feature selection a key factor in optimizing processor models and policy-learning algorithms.
2.3 Performance Unpredictability
Energy consumption in multi-core processors depends on multiple features, including frequency, voltage, temperature, and performance data from each core. As shown in Figure 1, energy consumption can vary by an order of magnitude when the same frequency and number of cores are used across different processors. This variability highlights the importance of feature selection in identifying the most significant predictors of energy consumption across different platforms, frequencies, and core configurations. Understanding which features retain their relevance across diverse conditions aids in developing a global policy for DVFS and task-to-core allocation.
Figure 1: Energy consumption variation across three different processors with identical frequency and core count settings: Intel Core i7 8th gen. (Corei78) with 4 cores, Intel Core i7 8th gen. (Corei712) with 14 cores, and Intel Xeon 2680 v3 (Xeon) with 12 cores. The results are shown for three different OpenMP benchmarks.
3 Design Methodology
Our objective is to develop a multi-stage methodology that selects the most critical features and applies these insights to optimize thermal behavior and energy efficiency in multi-core processors. In what follows, we present three feature selection approaches—filter-based, wrapper-based, and embedded—and show how their combined use informs an intelligent task-to-core allocation algorithm.
3.1 Filter-Based Feature Selection
Filter methods assess the relevance of features by examining intrinsic properties of the data without involving any learning algorithms. One common technique in filter methods is to evaluate the correlation between features and the target variable or among the features themselves. In our study, we utilize the Pearson correlation coefficient Sedgwick (2012) to quantify the linear relationship between core temperatures, which can indicate adjacency and potential heat transfer between cores.
Given a set of $n$ observations for $m$ cores, let $\theta_{i,k}$ denote the temperature of core $i$ at observation $k$, and $\bar{\theta}_{i}$ represent the mean temperature of core $i$ across all observations. The Pearson correlation coefficient $r_{ij}$ between cores $i$ and $j$ is computed as:
$r_{ij}=\frac{\sum_{k=1}^{n}\bigl{(}\theta_{i,k}-\bar{\theta}_{i}\bigr{)}\bigl{(}\theta_{j,k}-\bar{\theta}_{j}\bigr{)}}{\sqrt{\sum_{k=1}^{n}\bigl{(}\theta_{i,k}-\bar{\theta}_{i}\bigr{)}^{2}}\,\sqrt{\sum_{k=1}^{n}\bigl{(}\theta_{j,k}-\bar{\theta}_{j}\bigr{)}^{2}}}$ (1)
The value of $r_{ij}$ ranges from $-1$ to $1$, where $1$ indicates a perfect positive linear correlation, $-1$ indicates a perfect negative linear correlation, and 00 signifies no linear correlation between the temperatures of cores $i$ and $j$. A high positive correlation suggests that the temperatures of the two cores rise and fall together, possibly due to physical proximity and shared thermal characteristics.
By constructing a correlation matrix $\mathbf{R}=[r_{ij}]$ for all pairs of cores, we can visualize and identify clusters of cores that are thermally correlated. This information is crucial for designing task-to-core allocation strategies that minimize thermal hotspots. As shown in Figure 2, the lower diagonal part represents the regression line in the sparsified data, and the upper diagonal part shows the colored correlation matrix, where a greener color indicates a more positive correlation between the temperatures of two cores.
Figure 2: Correlation matrix based on Pearson correlation coefficients for 10 selected cores from an Intel Core i7 12th Gen processor with 14 cores.
Correlation-Aware Task-to-Core Allocation Algorithm.
To leverage the insights from the correlation analysis, we propose a correlation-aware task-to-core allocation algorithm. The goal is to assign tasks to cores that are less thermally correlated, thereby reducing the risk of localized overheating and improving overall energy efficiency.
Let $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$ denote the set of available cores, and let $\mathbf{R}$ be the correlation matrix computed using Equation (1). The algorithm proceeds as follows:
  1. 1.
    Compute Core Correlation Scores: For each core $c_{i}$, calculate a correlation score $s_{i}$ defined as the average absolute correlation between core $c_{i}$ and all other cores:
    $s_{i}=\frac{1}{m-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}|r_{ij}|$ (2)
    A lower score $s_{i}$ indicates that core $c_{i}$ is less correlated with other cores.
  2. 2.
    Rank Cores Based on Correlation Scores: Sort the cores in ascending order of their correlation scores to obtain a ranked list $\mathcal{C}_{\text{ranked}}$.
  3. 3.
    Select Cores for Task Assignment: Given the number of tasks $T$ to be assigned, select the first $T$ cores from $\mathcal{C}_{\text{ranked}}$, which are the least correlated cores.
  4. 4.
    Assign Tasks to Selected Cores: Allocate tasks to the selected cores, ensuring that each task is assigned to a core with minimal thermal correlation to other active cores.
  5. 5.
    Update Temperature Buffer: After task execution, update the temperature observations $\theta_{i,k}$ to reflect the new core temperatures, and recompute the correlation matrix $\mathbf{R}$ for subsequent allocations.
This algorithm dynamically adapts to the thermal behavior of the system, effectively distributing the thermal load more evenly across the cores. Figure 3 presents a comparison of average temperature and energy consumption between correlation-based and random core selection across different governors. While random core allocation is generally expected to yield better results due to its unbiased distribution, the correlation-based approach demonstrates comparable performance, even when allocating a small number of cores to each workload. In scenarios where the order of core allocation is crucial—such as selecting 10 out of 14 available cores—the correlation-aware allocation method proves to be more effective in reducing both energy consumption and temperature, offering an advantage over random allocation.
Figure 3: Comparison of average temperature and energy consumption for correlation-based (Corr) and random (Rand) core selection. Experiments performed on intel Core i7 12th Gen processor with 14 cores.
3.2 Wrapper-Based Feature Selection
Wrapper methods evaluate subsets of features using a predictive model. Because they account for feature interactions, they typically yield higher accuracy than filter methods in finding the importance of the features on a specific feature parameter but may be more computationally expensive.
We employ the backward stepwise selection algorithm, which starts with all available features and iteratively removes the least significant feature based on a specified criterion. In our case, we use the Ordinary Least Squares (OLS) regression model to predict the target variables (energy consumption and average temperature) and assess the significance of the characteristics using statistical tests.
Backward Stepwise Selection Algorithm.
Let =$\mathcal{F}=\{x_{1},x_{2},\ldots,x_{d}\}$ denote the full set of features. The backward stepwise selection algorithm proceeds as follows:
  1. 1.
    Initial Model: Fit the OLS regression model using all features in $\mathcal{F}$:
    $y=\beta_{0}+\sum_{i=1}^{d}\beta_{i}x_{i}+\varepsilon$ (3)
    where $y$ is the target variable, $\beta_{0}$ is the intercept, $\beta_{i}$ are the coefficients, and $\varepsilon$ is the error term.
  2. 2.
    Evaluate Feature Significance: For each feature $x_{i}$, compute the t-statistic and the corresponding p-value to assess its statistical significance. The t-statistic for coefficient $\beta_{i}$ is calculated as:
    $t_{i}=\frac{\hat{\beta}_{i}}{\text{SE}(\hat{\beta}_{i})},$ (4)
    where $\hat{\beta}_{i}$ is the estimated coefficient and $\text{SE}(\hat{\beta}_{i})$ is its standard error.
  3. 3.
    Feature Elimination: Identify the feature with the highest p-value (least significant) that exceeds a predefined significance level (e.g., $\alpha=0.05$). Remove this feature from the model.
  4. 4.
    Iterative Refinement: Refit the OLS model using the reduced feature set and repeat steps 2 and 3 until all remaining features are statistically significant.
  5. 5.
    Model Selection Criteria: At each iteration, evaluate the model using metrics such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), Mallows’ $C_{p}$, Adjusted $R^{2}$, and Cross-Validation Error (CV Error). These metrics help balance model complexity and goodness of fit.
Evaluation Metrics for Our Wrapper Algorithm.
We employ multiple metrics to gauge not only how well each model fits the data, but also how efficiently it uses the available features. This multi-criteria evaluation helps us to verify a model that performs reliably, avoids overfitting, and remains computationally viable for energy-aware and thermal-critical environments.
  • Akaike Information Criterion (AIC): AIC estimates the relative quality of statistical models for a given dataset:
    $\text{AIC}=2k-2\ln(L),$ (5)
    where $k$ is the number of estimated parameters, and $L$ is the maximized value of the likelihood function. Lower AIC values imply better trade-offs between model complexity and fit.
  • Bayesian Information Criterion (BIC): BIC imposes a stronger penalty on model complexity than AIC:
    $\text{BIC}=k\ln(n)-2\ln(L),$ (6)
    where $n$ is the number of observations. BIC is helpful for avoiding over-complex models in resource-constrained environments.
  • Mallows’ $C_{p}$: This criterion assesses the balance between the model’s complexity and its fit to the data:
    $C_{p}=\frac{\text{RSS}}{\hat{\sigma}^{2}}-(n-2k),$ (7)
    where RSS is the residual sum of squares, and $\hat{\sigma}^{2}$ is an estimate of the error variance.
  • Adjusted $R^{2}$: Adjusted $R^{2}$ modifies the coefficient of determination ($R^{2}$) to account for the number of predictors:
    Adjusted $\text{Adjusted }R^{2}=1-\left(\frac{(1-R^{2})(n-1)}{n-k-1}\right).$ (8)
    Unlike plain $R^{2}$, it penalizes the model for including uninformative features.
  • Cross-Validation Error: CV Error (often computed via $K$-fold CV) offers an unbiased measure of out-of-sample performance, revealing the model’s generalization capability and mitigating overfitting concerns.
By applying these evaluation metrics, we identify the optimal number of features that balance predictive accuracy and model simplicity. Figures 4(a) and 4(b) demonstrate that using fewer than 8 features suffices for accurate estimation of both average temperature and energy consumption, indicating that tracking only the most relevant predictors can improve energy efficiency and thermal behavior.
(a) Average temperature estimation
(b) Energy consumption estimation
(c) Determining the significance of features using Random Forest with respect to cross-validation error on Intel Core i7 12th Gen.
(d) Comparison of FCN models with and without feature selection and bootstrapping on Intel Core i7 12th Gen.
Figure 4: (Left) Backward stepwise selection for estimating energy consumption and average temperature. Retaining fewer than 8 predictors (features) yields accurate predictions in both cases. (Right) Feature importance analysis and comparison of FCN models with and without feature selection/bootstrapping. All experiments performed on Intel Core i7 12th Gen.
3.3 Embedded Feature Selection Using Random Forest
Embedded methods incorporate feature selection into the model training process, thereby minimizing the need for multiple model evaluations on different feature subsets. In this work, we employ a Random Forest (RF) regressor, an ensemble approach that constructs multiple decision trees and aggregates their predictions.
Random Forest Algorithm.
The RF algorithm proceeds as follows:
  1. 1.
    Bootstrap Sampling: Generate $N$ bootstrap samples from the original dataset.
  2. 2.
    Tree Construction: For each bootstrap sample, grow a regression tree by selecting a random subset of features at each node (often $\sqrt{d}$ features). Each tree is grown to its maximum depth without pruning, although hyperparameters such as the number of trees or maximum depth can be tuned for embedded constraints.
  3. 3.
    Prediction Aggregation: For regression tasks, the final prediction is the average of the individual tree outputs:
    $\hat{y}=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}.$ (9)
Feature Importance Computation.
Random Forests provide a measure of the importance of features by evaluating the total decrease in the impurity of the nodes in all trees. For regression trees, impurity is commonly measured by the residual sum of squares. The importance score $I_{j}$ for feature $x_{j}$ is thus:
$I_{j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t\in T_{i}}\Delta I_{t,j},$ (10)
where $\Delta I_{t,j}$ is the impurity decrease at node $t$ when splitting on $x_{j}$, and $T_{i}$ is the set of nodes in the $i$-th tree. The ranking of features by these importance scores guides the selection of highly influential predictors.
Bootstrapping for Data Augmentation.
To increase model robustness, we employ bootstrapping, a resampling method that draws multiple datasets of size $n$ with replacement. Let $\mathcal{D}$ be the original dataset of size $n$. Forming $B$ bootstrap samples $\{\mathcal{D}_{1},\ldots,\mathcal{D}_{B}\}$ helps estimate variance and stabilize the final model through aggregation of predictions.
Environment Modeling with Feature Selection.
After identifying the most significant features using RF importance scores, we build predictive models for environment modeling—particularly neural networks—tailored to embedded constraints. Let $\mathbf{x}\in\mathbb{R}^{p}$ (with $p<d$) denote the reduced feature set. We train a Fully Connected Neural Network (FCN) with multiple layers and non-linear activations to predict key variables such as energy consumption or temperature. The network is trained by minimizing the mean squared error (MSE):
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\bigl{(}y_{i}-f(\mathbf{x}_{i};\theta)\bigr{)}^{2},$ (11)
where $f(\mathbf{x}_{i};\theta)$ is the network’s output for input $\mathbf{x}_{i}$ with parameters $\theta$, and $y_{i}$ is the true label. By restricting the input to a smaller set of highly relevant features, the FCN requires fewer parameters and less computation, making it feasible for real-time applications.
3.4 Summary and Method Synergy
We combine:
  • Filter-Based Selection (Pearson correlation) to reveal thermally correlated cores, guiding an algorithm for task-to-core allocation.
  • Wrapper-Based Selection (Backward Stepwise) to refine features for single-target regression (energy consumption or temperature) with high accuracy.
  • Embedded Selection (Random Forest) to account for non-linear relationships, automatically ranking features and integrating bootstrapping to reduce variance.
These complementary methods yield a multi-stage strategy: We first detect obvious correlations, then prune unnecessary predictors using a regression-based wrapper, and finally leverage a Random Forest to capture residual non-linearities. The ultimate outcome is a set of crucial features that guides scheduling decisions and real-time predictive modeling, simultaneously optimizing thermal distribution and energy usage in multi-core processors.
4 Experiments
This section describes the experimental setup, including platforms, benchmarks, and implementation details, followed by an in-depth discussion of results that highlight the effectiveness of our feature selection and modeling strategies.
4.1 Experimental Platform, Benchmarks, and Evaluation
The results and figures in this paper are extracted from superscalar Intel Core i7 12th Gen with 14 cores but also verified on Intel core i7 8th, Intel Xeon 2680, and Jetson TX2. Each platform was evaluated on three metrics: makespan, energy consumption, and average core temperature to give importance to the features. All experiments targeted the Barcelona OpenMP Tasks (BOTs) suite Duran et al. (2009), which provides diverse parallel workloads.
4.2 Implementation and Training Details
Before model training, data from each platform’s profiler and temperature sensors were split into training and test sets. Key hyperparameters included batch size, number of hidden neurons, epoch count, and learning rate, with Mean Squared Error (MSE) as the primary loss criterion. We explored various neural network architectures—Fully Connected Networks (FCN), Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM) networks, Convolutional Neural Networks (CNN), and Attention-based models (e.g., Transformers)—and retained the best-performing variants based on validation or training loss.
We derived feature subsets using Pearson correlation, backward stepwise OLS selection, or random-forest-based importance rankings. To reduce overfitting and increase robustness, a bootstrapping approach was adopted, resampling the selected features to create multiple training sets. This not only preserved critical predictors for energy and temperature estimation but also offered insight into how these features interacted in different sampling scenarios.
4.3 Empirical Results
Two predictive models were developed: one for profiler output (energy consumption, context switches, branch misses, etc.) and another for future temperature values. In the profiler model, we used every available profiling feature on each platform (e.g., cache miss rates, branch miss rates, CPU cycles, instructions per cycle, average speed, page faults), together with current per-core temperatures $\theta_{i}$ and temperature differences $\Delta\theta_{i}$. Conversely, the temperature model was limited to each core’s temperature and its difference over time. For the Intel Core i7 12th Gen with 14 cores, we specifically included all 14 temperature readings, their differences, and the average temperature to predict future thermal behavior. By selecting 30 features for the temperature model and 75 for the profiler model, we maintained enough diversity while avoiding unnecessary overhead.
Figure 4(c) shows how random forest-based feature selection and bootstrapping significantly lowered MSE in both models. Concentrating on fewer yet influential features reduced computational costs and often improved accuracy, an essential benefit for real-time embedded systems with strict power and thermal budgets.
During training, hyperparameters (e.g., learning rate, batch size) were tuned to balance convergence speed and generalization. Models were saved at their optimal checkpoint, identified when improvement plateaued or validation error declined notably. Independent neural networks were then trained and evaluated on temperature and profiler data, ensuring each specialized model addressed its corresponding objective effectively.
4.4 Comparative Model Analysis
We tested multiple neural network architectures to find the best design for energy and temperature prediction. Table 1 reports MSE values and parameter counts for the primary models and a state-of-the-art (SOTA) approach Hosseinimotlagh et al. (2021). Although baseline FCN models performed reasonably well, FCN variants with random forest feature selection (FCN+RF) surpassed the baseline in accuracy and parameter efficiency. Moreover, incorporating bootstrapping (FCN+RF+BS) further enhanced these metrics, yielding the smallest parameter counts and lowest MSE for both temperature and profiler tasks.
Table 1: MSE and total number of parameters for different architectures.
Model Temperature Profiler
MSE Params MSE Params
FCN 1.0299 2014 3.9047 3787
FCN+RF 0.9808 1694 2.4862 2699
FCN+RF+BS 0.9640 1694 2.4669 2699
RNN 1.0119 3070 2.8493 4843
LSTM 1.0307 9310 2.7778 15115
Conv 1.0134 5118 2.8217 6891
Attention 1.0143 6238 2.8933 8011
SOTA 2.5000 - - -
Figure 4(d) illustrates the effect of feature selection and bootstrapping on test MSE. Omitting either often resulted in inferior performance, emphasizing the benefit of pruning superfluous predictors and employing resampling. By centering on a streamlined feature subset, the final FCN models boosted prediction accuracy and remained computationally light, confirming the multi-stage selection and resampling methodology.
Overall, combining advanced feature selection (filter, wrapper, and embedded) with neural network models substantially improved prediction for energy and temperature, meeting real-time and power constraints in multi-core embedded systems. The synergy among reduced feature sets, resampling techniques, and specialized architectures (e.g., FCN+RF+BS) proved effective for these workloads.
5 Related Work
Probabilistic Methods for DVFS and Task-to-Core Allocation.
Estimating probabilistic worst-case execution time (pWCET) and worst-case energy consumption (pWCEC) in embedded real-time systems has been explored via measurement-based and static approaches Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017). Pallister et al. Pallister et al. (2017) analyzed the impact of different instructions on pWCEC, while the extreme value theorem (EVT) Edgar and Burns (2001) predicted upper bounds on performance metrics. However, most existing probabilistic models do not systematically evaluate the statistical importance of system and performance metrics for bounding energy or latency. Our work introduces a statistical learning approach that prioritizes feature significance, enabling low-energy designs and enhanced environment modeling under DVFS and task-to-core allocation.
Statistical Learning.
Researchers have applied statistical learning to identify influential features for energy optimization and scheduling, often based on hardware events or application parameters Sasaki et al. (2007); Cazorla et al. (2019); Liu et al. (2021). Sasaki et al. Sasaki et al. (2007) used decision trees to minimize table look-up overhead in DVFS settings, and Cazorla et al. Cazorla et al. (2019) examined hardware counters to reduce energy consumption. Liu et al. Liu et al. (2021) ranked compiler-generated features by correlation with latency. However, these studies overlooked the potential of runtime performance metrics, sampling techniques, and accuracy trade-offs. We address these gaps by comprehensively evaluating feature correlations and devising a feature selection strategy suitable for parallel scheduling and environment modeling.
Low-Energy DVFS and Task-to-Core Allocation.
A broad body of work investigates low-energy scheduling on multicore platforms using DVFS and task-to-core assignment Xie et al. (2021, 2017); Zhu et al. (2004); Jiang et al. (2019); Chen et al. (2018); Zhou et al. (2018); Kim and Wu (2020); Dinakarrao et al. (2019); Shen et al. (2012); Wang et al. (2017). Xie et al. Xie et al. (2021) reviewed heuristic, meta-heuristic, and machine learning algorithms for parallel scheduling under energy constraints. Despite these advances, many machine learning methods for low-energy DVFS management demand large datasets and incur high computational overhead. Our statistical learning approach alleviates these drawbacks by reducing data requirements while preserving prediction accuracy, making it practical for embedded and heterogeneous processing applications.
Few-Shot RL.
Few-shot learning techniques—including transfer learning, meta-learning, and data augmentation—target reduced data collection in processor scheduling Wang et al. (2020); Lee et al. (2020); Wang et al. (2016); Florensa et al. (2017); Arora and Doshi (2021). Model-agnostic meta-learning (MAML) Finn et al. (2017) adapts to new tasks with minimal data, and model-based reinforcement learning (RL) Moerland et al. (2023) approximates transition functions to reduce reliance on real-world samples. Prior studies Lin et al. (2023); Kim et al. (2021); Zhou and Lin (2021); Zhang et al. (2024) have not fully explored statistical resampling or feature importance evaluations for energy-efficient task-to-core scheduling. Our approach integrates these elements, bridging the gap by combining statistical resampling with a robust feature selection methodology.
Feature Evaluation.
Predictive models for thermal-aware scheduling have been proposed in numerous works Yan et al. (2003); Brooks et al. (2007); Maity et al. (2022); Hosseinimotlagh and Kim (2019); Singla et al. (2015); Li and Wu (2012); Kassab et al. (2021), often relying on utilization or temperature monitoring to avoid thermal throttling Maity et al. (2022); Lin et al. (2023); Kim et al. (2021). Although these models forecast thermal behavior based on transient and ambient states Hosseinimotlagh et al. (2021), they lack a rigorous statistical correlation analysis for feature selection. We address this limitation with a systematic approach that examines feature interdependencies and their effects on energy efficiency and performance, thereby optimizing DVFS strategies with more precise task-to-core allocation.
6 Conclusion
We demonstrated the effectiveness of feature selection using statistical learning for environment modeling and task-to-core allocation in embedded systems. Our correlation-aware task-to-core allocation reduces energy consumption by up to 10% and temperature by up to 5C compared to random core selection. The compressed bootstrapped regression model reduces thermal prediction error by 6% and the number of parameters by 16%. Tested on Intel Core i7 8th and 12th generation, Intel Xeon 2680 processors and Jetson TX2, our method shows a 61.6% reduction in mean squared error compared to state-of-the-art approach. This finding paves the way for future use of statistical learning methods in performance efficiency of task-to-core allocation in heterogeneous processors.