### v2

# Feature-Aware Task-to-Core Allocation in Embedded Multi-core Platforms via Statistical Learning

Abstract
Optimizing task-to-core allocation can substantially reduce power consumption in multi-core platforms without degrading user experience. However, existing approaches overlook critical factors such as parallelism, compute intensity, and heterogeneous core types. In this paper, we introduce a statistical learning approach for feature selection that identifies the most influential features—such as core type, speed, temperature, and application-level parallelism or memory intensity—for accurate environment modeling and efficient energy minimization, a critical consideration for embedded systems. Our experiments, conducted with state-of-the-art Linux governors and thermal modeling techniques, show that correlation-aware task-to-core allocation lowers energy consumption by up to 10% and reduces core temperature by up to 5C compared to random core selection. Furthermore, our compressed, bootstrapped regression model improves thermal prediction accuracy by 6% while cutting model parameters by 16%, yielding an overall mean square error reduction of 61.6% relative to existing approaches. We provided results based on superscalar Intel Core i7 12th Gen processors with 14 cores, and validated our method across a diverse set of hardware platforms and effectively balanced performance, power, and thermal demands through statistical feature evaluation.
Keywords Energy Optimization $\cdot$ Embedded Systems $\cdot$ Heterogeneous Platforms $\cdot$ Statistical Feature Evaluation $\cdot$ Task-to-Core Allocation
1 Introduction
Task-to-core allocation is a pivotal technique for enhancing the performance and reliability of embedded systems, where workloads exhibit distinct thermal and power consumption behaviors depending on core assignments. In multi-core processors, allocations can involve high-performance cores, low-energy cores, and graphic processing units (GPUs), each with unique performance characteristics across various configurations. Improper allocation can lead to thermal overheating, triggering throttling and reducing chip reliability and lifespan Hosseinimotlagh and Kim (2019). Cooling costs for overheated chips are significant, approximately \$3 per watt of heat dissipation Skadron et al. (2003). While dynamic voltage and frequency scaling (DVFS) complements allocation by adjusting power dynamically—potentially reducing consumption by 75% without altering user experience Ratković et al. (2015)—our focus is on optimizing task-to-core allocation to manage thermal and energy constraints in embedded systems. Existing approaches often rely on historical workload and sensor data to predict future behavior Maity et al. (2022), but lack a systematic, statistically robust framework, which our hybrid methodology addresses with novelty and precision.
Embedded systems, foundational to automotive, telecommunications, and consumer electronics, must operate under strict power and thermal constraints while delivering high performance. Task-to-core allocation is crucial for managing these constraints in heterogeneous architectures with diverse core types (e.g., high-performance, low-power, GPUs). Feature selection and evaluation identify critical features and correlations, enabling optimal core assignments for thermal stability and energy efficiency Shekarisaz et al. (2021). For instance, in real-time automotive systems, allocation ensures compute-intensive tasks run on performance cores without overheating, while mobile devices assign power-hungry applications to high-performance cores and background tasks to low-power cores, extending battery life. Identifying features tied to power and thermal behavior—such as temperature, frequency, and core adjacency—establishes general rules for allocation, enhancing embedded system reliability. This paper introduces a hybrid statistical learning framework, integrating Random Forest (RF), backward stepwise selection, and correlation analysis, to optimize task-to-core allocation across platforms, prioritizing allocation over DVFS for embedded contexts.
Application performance depends on execution factors like core type, speed, temperature, and application characteristics—memory or compute intensiveness and parallelism level Pagani et al. (2018). Linux governors Brodowski et al. (2013) and related studies often focus on CPU utilization for DVFS policies, targeting latency, temperature, or energy Liu et al. (2021); Lin et al. (2023); Kim et al. (2021). However, for task-to-core allocation, characteristics like parallelism, memory/compute intensiveness, and branch counts are equally critical. Applications with frequent branch misses benefit from sequential execution on single cores, while parallel, compute-intensive tasks excel on GPUs or performance cores in heterogeneous platforms. Moreover, temperature correlations between cores indicate adjacency and overheating risks. Our approach transcends utilization-centric methods, employing a multi-stage statistical framework to capture a comprehensive feature set and their interdependencies, ensuring clarity and applicability in embedded systems.
Collecting real data post-execution for task-to-core allocation optimization is computationally expensive and imprecise due to hardware sampling delays. Instead, allocation decisions can leverage inference from trained environmental models that account for randomness and feature importance to mitigate overfitting Liu et al. (2021). An augmentation step further reduces sampling overhead, boosting efficiency. Complementary approaches include hierarchical multi-agent DVFS scheduling Pivezhandi et al. (2026b), zero-shot LLM-guided allocation Pivezhandi and Saifullah (2026b), flow-augmented few-shot RL for sensor data Pivezhandi and Saifullah (2026a), and graph-driven performance modeling Pivezhandi et al. (2026a). Our paper introduces the first statistical learning framework for embedded systems, uniquely integrating RF-based feature reduction, backward stepwise selection, and correlation-aware allocation with bootstrapping to create a robust, platform-agnostic model. Unlike prior heuristic or filter-based methods, which lack systematic feature evaluation or adaptability across heterogeneous platforms, our approach outperforms existing techniques—where they exist—achieving up to 10% energy savings and 61.6% lower mean squared error, as validated against SOTA baselines on diverse embedded hardware.
Little work applies statistical learning to feature selection for task-to-core allocation with the rigor we propose. Statistical learning predicts outcomes and reveals feature relationships. Prior efforts used extreme value theorem (EVT) for energy and execution time bounds in DVFS contexts Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017), while Liu et al. Liu et al. (2021) applied XGBoost for latency-related features, and Sasaki et al. Sasaki et al. (2007) used decision trees and correlation for energy per instruction. These lack the integrated, multi-method focus we introduce for allocation, synergizing RF, OLS, and correlation analysis for superior predictive power and generalizability in embedded systems.
Our framework shows that correlation-aware task-to-core allocation reduces chip temperature at intermediate utilization by leveraging thermally independent cores. Backward stepwise selection reveals a small feature subset suffices for accurate energy estimation, while RF extracts critical features with low overhead. Bootstrapping enhances dataset robustness, yielding a compact, efficient model. This hybrid methodology surpasses prior single-method approaches, validated across embedded platforms for practical impact.
The key contributions are:
  1. 1.
    This paper pioneers a hybrid statistical learning approach, integrating RF, backward stepwise selection, and correlation analysis, to evaluate feature importance for accurate environment modeling and compressed learning in task-to-core allocation for embedded systems.
  2. 2.
    We implement correlation-aware allocation, statistically reducing temperature and energy by assigning tasks to uncorrelated cores, validated across heterogeneous embedded platforms.
  3. 3.
    We employ bootstrapping to augment data and boost accuracy, supporting few-shot and meta-learning in resource-constrained embedded settings.
  4. 4.
    We rigorously test against state-of-the-art methods, achieving up to 10% energy reduction and 5°C temperature decrease. Our compressed, bootstrapped model cuts prediction error by 6%, reduces parameters by 16%, and improves mean squared error by 61.6% over SOTA. Data were extracted from Intel Core i7 8th and 12th Gen and Intel Xeon 2680 processors.
In the remainder, Section 5 surveys statistical learning and scheduling, Section 2 details motivation and challenges, and Section 3 describes the methodology. Section 4 evaluates three processors, followed by conclusions in Section 6.
2 Motivation and Challenges
Feature selection is a cornerstone of statistical learning, pivotal for optimizing task-to-core allocation in embedded multi-core systems. It addresses critical challenges—curse of dimensionality, large data management, and performance unpredictability—common to platforms like Intel Core i7 and NVIDIA Jetson TX2. To surpass prior filter-based approaches and establish novelty, we leverage a hybrid methodology integrating Random Forest (RF), backward stepwise selection, and correlation analysis, tailored to embedded constraints. This section clarifies these challenges in the context of task-to-core allocation, emphasizing generalizability across heterogeneous architectures and grounding our motivation in rigorous, platform-validated insights.
2.1 Curse of Dimensionality
Feature selection is essential to mitigate the curse of dimensionality, a challenge amplified in embedded multi-core systems. As feature count rises—encompassing frequency, temperature, and performance metrics like cache misses—the data space grows exponentially, leading to sparsity in high-dimensional spaces Li et al. (2017). This sparsity risks model overfitting, degrading predictive performance on unseen data. In heterogeneous embedded systems, the diversity of cores (performance, low-energy, GPUs) expands the state space, with each core type contributing unique thermal and performance characteristics. For instance, Intel Core i7 12th Gen’s hybrid P/E-cores and Jetson TX2’s big-LITTLE clusters introduce complex feature interactions. Efficient feature selection, as implemented via our RF-based reduction, reduces this complexity, enabling compact, generalizable models critical for real-time embedded applications—offering a novel advance over traditional dimensionality handling.
2.2 Large Data Management
Managing the growing volume of data is a pressing challenge for task-to-core allocation in embedded systems, where resource constraints demand efficiency. Allocation strategies may use static historical data or streaming inputs adapting to runtime conditions. With streaming data’s rise—common in automotive or mobile embedded platforms—memory management becomes critical, as unpredictable data volumes can overwhelm limited resources. Retaining unnecessary features, such as redundant performance counters, inflates storage and processing overhead, straining embedded systems. Our hybrid approach, combining RF for initial feature pruning and bootstrapping for data augmentation, optimizes processor models and allocation algorithms, reducing overhead while maintaining accuracy. This addresses large data challenges with clarity and applicability across static and dynamic embedded environments, surpassing simpler filter-based data handling.
2.3 Performance Unpredictability
Performance unpredictability in embedded multi-core processors stems from energy and thermal variability, driven by features like temperature, frequency, and core-specific performance data. Figure 2 illustrates this, showing energy consumption varying by an order of magnitude across identical frequency and core settings on Intel Core i7 8th Gen (4 cores), Intel Core i7 12th Gen (14 cores), and Intel Xeon 2680 v3 (12 cores) under OpenMP benchmarks. This variability underscores the need for feature selection to pinpoint significant predictors of energy and thermal behavior across platforms and configurations. Our methodology—using backward stepwise OLS for energy-critical features and correlation analysis for thermal independence—identifies robust predictors, enabling a global allocation policy. Unlike DVFS-centric prior works, we prioritize task-to-core allocation, validated rigorously with up to 10% energy savings and 5°C temperature reductions, ensuring practical impact in embedded systems.
Figure 1: Energy consumption variation across three different processors with identical frequency and core count settings: Intel Core i7 8th Gen (Corei78) with 4 cores, Intel Core i7 12th Gen (Corei712) with 14 cores, and Intel Xeon 2680 v3 (Xeon) with 12 cores. The results are shown for three different OpenMP benchmarks.
Figure 2: Determining the significance of features using Random Forest with respect to cross-validation error on Intel Core i7 12th Gen.
3 Design Methodology
Our objective is to develop a robust multistage methodology based on three feature selection approaches—embedded, wrapper-based, and filter-based—and show how their combined use informs an intelligent task-to-core allocation algorithm for embedded multi-core systems. We introduce a hybrid framework integrating Random Forest (RF), backward stepwise OLS, and Pearson correlation, surpassing traditional filter-based methods in efficiency and predictive power. We provide a structured workflow adaptable to diverse platforms like Intel Core i7 8th and 12th Gen, Intel Xeon 2680 v3, and NVIDIA Jetson TX2, while ensuring experimental rigor through validated outcomes (e.g., 10% energy savings, 5°C temperature reduction). This section details each stage, linking to embedded system optimization.
3.1 Data Collection and Environment Setup
We begin by profiling the target embedded hardware under diverse workloads to capture comprehensive data on energy consumption, temperature variations, and relevant performance metrics. The granularity of thermal sensors significantly varies across platforms. Intel Core i7 8th generation and Intel Xeon 2680 v3 processors provide homogeneous core architectures with individual per-core thermal sensors, enabling detailed per-core correlation analyses. In contrast, the Intel Core i7 12th generation features a heterogeneous architecture composed of performance (P-cores) and efficient (E-cores) cores, each type operating at different frequencies and performance characteristics. Similarly, the NVIDIA Jetson TX2 offers thermal readings aggregated at the cluster level, typically distinguishing between big and little core clusters due to its big-LITTLE architecture. Thus, our methodology accommodates per-core, heterogeneous core, and cluster-level sensor capabilities, specifically tailored to embedded computing environments, ensuring broad applicability.
Our data collection covers a wide range of configurations, including varying core counts (4, 12, or 14 cores), diverse CPU frequencies, and task prioritizations. Tasks include representative benchmarks such as Fibonacci (fib), matrix multiplication (Strassen), and sequence alignment (alignment), providing varied computational intensities and memory access patterns typical of embedded system workloads. This diverse dataset enables accurate modeling of core behaviors under different workloads and thermal conditions, essential for embedded system optimization.
Initially, the system setup involves configuring the frequency governor (typically schedutil) and initializing a temperature buffer capable of storing historical temperature data from cores (up to 10,000 entries). Additionally, a detailed task history is maintained, capturing execution parameters such as energy consumption, profiler metrics, and temperature data under varying task priorities and frequencies. This systematic variation ensures the dataset’s robustness and representativeness for embedded systems.
Energy monitoring employs built-in hardware interfaces such as /sys/class/powercap/intel-rapl/ on Intel processors or on-board power sensors for the Jetson TX2, recording total energy consumed or average power during task execution. Performance metrics collected include CPU frequency, utilization, memory bandwidth, and hardware performance counters like cache misses, gathered via utilities such as perf and cpufreq-info at consistent intervals.
The random core assignment strategy serves as a baseline to evaluate the effectiveness of our proposed correlation-aware core allocation strategy. In the random assignment approach, tasks run concurrently on randomly selected cores, typically half of the available cores excluding core 0, reserved for system tasks. Post-execution, we record metrics including temperature, energy consumption, and performance data. Execution is temporarily paused if core temperatures exceed predefined thresholds (e.g., depending on thermal throttling point of Intel Core i7 which is 100°C), allowing cores to cool down before resuming execution. This baseline facilitates comparative analysis against the correlation-aware allocation strategy.
The collected dataset serves as input for the subsequent embedded feature reduction process, employing Random Forest algorithms to identify minimal yet critical feature subsets. This refined feature set is then utilized for accurate environment modeling, guiding intelligent, temperature-correlation-driven task-to-core allocation strategies specifically designed for embedded systems. Ultimately, our structured, multi-stage methodology aims to optimize thermal management and energy efficiency across diverse embedded multi-core computing platforms.
3.2 Embedded Feature Selection Using Random Forest
Embedded methods incorporate feature selection into the model training process, thereby minimizing the need for multiple model evaluations on different feature subsets. In this work, we employ a Random Forest (RF) regressor, an ensemble approach that constructs multiple decision trees and aggregates their predictions, offering a novel low-overhead alternative to filter-based methods.
Random Forest Algorithm.
The RF algorithm proceeds as follows:
  1. 1.
    Bootstrap Sampling: Generate $N$ bootstrap samples from the original dataset.
  2. 2.
    Tree Construction: For each bootstrap sample, grow a regression tree by selecting a random subset of features at each node (often $\sqrt{d}$ features). Each tree is grown to its maximum depth without pruning, although hyperparameters such as the number of trees or maximum depth can be tuned for embedded constraints.
  3. 3.
    Prediction Aggregation: For regression tasks, the final prediction is the average of the individual tree outputs:
    $\hat{y}=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}.$ (1)
Feature Importance Computation.
Random Forests provide a measure of the importance of features by evaluating the total decrease in the impurity of the nodes in all trees. For regression trees, impurity is commonly measured by the residual sum of squares. The importance score $I_{j}$ for feature $x_{j}$ is thus:
$I_{j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t\in T_{i}}\Delta I_{t,j},$ (2)
where $\Delta I_{t,j}$ is the impurity decrease at node $t$ when splitting on $x_{j}$, and $T_{i}$ is the set of nodes in the $i$-th tree. The ranking of features by these importance scores guides the selection of highly influential predictors.
Bootstrapping for Data Augmentation.
To increase model robustness, we employ bootstrapping, a resampling method that draws multiple datasets of size $n$ with replacement. Let $\mathcal{D}$ be the original dataset of size $n$. Forming $B$ bootstrap samples $\{\mathcal{D}_{1},\ldots,\mathcal{D}_{B}\}$ helps estimate variance and stabilize the final model through aggregation of predictions.
Environment Modeling with Feature Selection.
After identifying the most significant features using RF importance scores, we build predictive models for environment modeling—particularly neural networks—tailored to embedded constraints. Let $\mathbf{x}\in\mathbb{R}^{p}$ (with $p<d$) denote the reduced feature set. We train a Fully Connected Neural Network (FCN) with multiple layers and non-linear activations to predict key variables such as energy consumption or temperature. The network is trained by minimizing the mean squared error (MSE):
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\bigl(y_{i}-f(\mathbf{x}_{i};\theta)\bigr)^{2},$ (3)
where $f(\mathbf{x}_{i};\theta)$ is the network’s output for input $\mathbf{x}_{i}$ with parameters $\theta$, and $y_{i}$ is the true label. By restricting the input to a smaller set of highly relevant features, the FCN requires fewer parameters and less computation, making it feasible for real-time applications. As shown in Figure 2, our results indicate that using 39 out of 72 predictors provides nearly the same error value for estimating the future state of the processor profiler data model. For the sensor temperature model, using 21 out of 31 features yields the optimal Cross-Validation error (CV error), validated for embedded efficiency.
3.3 Wrapper-Based Feature Selection
Wrapper methods evaluate subsets of features using a predictive model. Because they account for feature interactions, they typically yield higher accuracy than filter methods in finding the importance of the features on a specific feature parameter but may be more computationally expensive. We employ the backward stepwise selection algorithm, which starts with all available features and iteratively removes the least significant feature based on a specified criterion. In our case, we use the Ordinary Least Squares (OLS) regression model to predict the target variables (energy consumption and average temperature) and assess the significance of the characteristics using statistical tests, enhancing the hybrid framework’s precision.
Backward Stepwise Selection Algorithm.
Let $\mathcal{F}=\{x_{1},x_{2},\ldots,x_{d}\}$ denote the full set of features. The backward stepwise selection algorithm proceeds as follows:
  1. 1.
    Initial Model: Fit the OLS regression model using all features in $\mathcal{F}$:
    $y=\beta_{0}+\sum_{i=1}^{d}\beta_{i}x_{i}+\varepsilon$ (4)
    where $y$ is the target variable, $\beta_{0}$ is the intercept, $\beta_{i}$ are the coefficients, and $\varepsilon$ is the error term.
  2. 2.
    Evaluate Feature Significance: For each feature $x_{i}$, compute the t-statistic and the corresponding p-value to assess its statistical significance. The t-statistic for coefficient $\beta_{i}$ is calculated as:
    $t_{i}=\frac{\hat{\beta}_{i}}{\text{SE}(\hat{\beta}_{i})},$ (5)
    where $\hat{\beta}_{i}$ is the estimated coefficient and $\text{SE}(\hat{\beta}_{i})$ is its standard error.
  3. 3.
    Feature Elimination: Identify the feature with the highest p-value (least significant) that exceeds a predefined significance level (e.g., $\alpha=0.05$). Remove this feature from the model.
  4. 4.
    Iterative Refinement: Refit the OLS model using the reduced feature set and repeat steps 2 and 3 until all remaining features are statistically significant.
  5. 5.
    Model Selection Criteria: At each iteration, evaluate the model using metrics such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), Mallows’ $C_{p}$, Adjusted $R^{2}$, and CV Error. These metrics help balance model complexity and goodness of fit.
Evaluation Metrics for Our Wrapper Algorithm.
We employ multiple metrics to gauge not only how well each model fits the data, but also how efficiently it uses the available features. This multi-criteria evaluation helps us verify a model that performs reliably, avoids overfitting, and remains computationally viable for energy-aware and thermal-critical environments:
  • Akaike Information Criterion (AIC): AIC estimates the relative quality of statistical models for a given dataset:
    $\text{AIC}=2k-2\ln(L),$ (6)
    where $k$ is the number of estimated parameters, and $L$ is the maximized value of the likelihood function. Lower AIC values imply better trade-offs between model complexity and fit.
  • Bayesian Information Criterion (BIC): BIC imposes a stronger penalty on model complexity than AIC:
    $\text{BIC}=k\ln(n)-2\ln(L),$ (7)
    where $n$ is the number of observations. BIC is helpful for avoiding over-complex models in resource-constrained environments.
  • Mallows’ $C_{p}$: This criterion assesses the balance between the model’s complexity and its fit to the data:
    $C_{p}=\frac{\text{RSS}}{\hat{\sigma}^{2}}-(n-2k),$ (8)
    where RSS is the residual sum of squares, and $\hat{\sigma}^{2}$ is an estimate of the error variance.
  • Adjusted $R^{2}$: Adjusted $R^{2}$ modifies the coefficient of determination ($R^{2}$) to account for the number of predictors:
    $\text{Adjusted }R^{2}=1-\left(\frac{(1-R^{2})(n-1)}{n-k-1}\right).$ (9)
    Unlike plain $R^{2}$, it penalizes the model for including uninformative features.
  • Cross-Validation Error: CV Error (often computed via $K$-fold CV) offers an unbiased measure of out-of-sample performance, revealing the model’s generalization capability and mitigating overfitting concerns.
By applying these evaluation metrics, we identify the optimal number of features that balance predictive accuracy and model simplicity. Figures 3(a) and 3(b) demonstrate that using fewer than 8 features suffices for accurate estimation of both average temperature and energy consumption, indicating that tracking only the most relevant predictors can improve energy efficiency and thermal behavior, validated on Intel Core i7 12th Gen.
(a) Average temperature estimation
(b) Energy consumption estimation
Figure 3: Backward stepwise selection for estimating energy consumption and average temperature. Retaining fewer than 8 predictors (features) yields accurate predictions in both cases. Experiments performed on Intel Core i7 12th Gen.
Figure 4: Correlation matrix based on Pearson correlation coefficients for 10 selected cores from an Intel Core i7 12th Gen processor with 14 cores.
3.4 Filter-Based Feature Selection
Filter methods assess the relevance of features by examining intrinsic properties of the data without involving any learning algorithms. One common technique in filter methods is to evaluate the correlation between features and the target variable or among the features themselves. In our study, we utilize the Pearson correlation coefficient Sedgwick (2012) to quantify the linear relationship between core temperatures, which can indicate adjacency and potential heat transfer between cores, enhancing allocation decisions.
Given a set of $n$ observations for $m$ cores, let $\theta_{i,k}$ denote the temperature of core $i$ at observation $k$, and $\bar{\theta}_{i}$ represent the mean temperature of core $i$ across all observations. The Pearson correlation coefficient $r_{ij}$ between cores $i$ and $j$ is computed as:
$r_{ij}=\frac{\sum_{k=1}^{n}\bigl(\theta_{i,k}-\bar{\theta}_{i}\bigr)\bigl(\theta_{j,k}-\bar{\theta}_{j}\bigr)}{\sqrt{\sum_{k=1}^{n}\bigl(\theta_{i,k}-\bar{\theta}_{i}\bigr)^{2}}\,\sqrt{\sum_{k=1}^{n}\bigl(\theta_{j,k}-\bar{\theta}_{j}\bigr)^{2}}}$ (10)
The value of $r_{ij}$ ranges from 1-1 to $1$, where $1$ indicates a perfect positive linear correlation, 1-1 indicates a perfect negative linear correlation, and 0 signifies no linear correlation between the temperatures of cores $i$ and $j$. A high positive correlation suggests that the temperatures of the two cores rise and fall together, possibly due to physical proximity and shared thermal characteristics.
By constructing a correlation matrix $\mathbf{R}=[r_{ij}]$ for all pairs of cores, we can visualize and identify clusters of cores that are thermally correlated. This information is crucial for designing task-to-core allocation strategies that minimize thermal hotspots. As shown in Figure 4, the lower diagonal part represents the regression line in the sparsified data, and the upper diagonal part shows the colored correlation matrix, where a greener color indicates a more positive correlation between the temperatures of two cores.
Correlation-Aware Task-to-Core Allocation Algorithm.
To leverage the insights from the correlation analysis, we propose a correlation-aware task-to-core allocation algorithm. The goal is to assign tasks to cores that are less thermally correlated, thereby reducing the risk of localized overheating and improving overall energy efficiency.
Let $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$ denote the set of available cores, and let $\mathbf{R}$ be the correlation matrix computed using Equation (10). The algorithm proceeds as follows:
  1. 1.
    Compute Core Correlation Scores: For each core $c_{i}$, calculate a correlation score $s_{i}$ defined as the average absolute correlation between core $c_{i}$ and all other cores:
    $s_{i}=\frac{1}{m-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}|r_{ij}|$ (11)
    A lower score $s_{i}$ indicates that core $c_{i}$ is less correlated with other cores.
  2. 2.
    Rank Cores Based on Correlation Scores: Sort the cores in ascending order of their correlation scores to obtain a ranked list $\mathcal{C}_{\text{ranked}}$.
  3. 3.
    Select Cores for Task Assignment: Given the number of tasks $T$ to be assigned, select the first $T$ cores from $\mathcal{C}_{\text{ranked}}$, which are the least correlated cores.
  4. 4.
    Assign Tasks to Selected Cores: Allocate tasks to the selected cores, ensuring that each task is assigned to a core with minimal thermal correlation to other active cores.
  5. 5.
    Update Temperature Buffer: After task execution, update the temperature observations $\theta_{i,k}$ to reflect the new core temperatures, and recompute the correlation matrix $\mathbf{R}$ for subsequent allocations.
3.5 Summary of the Multi-Stage Methodology
  1. 1.
    Data Collection: Profile per-core or per-cluster temperatures, energy, and performance under varied workloads.
  2. 2.
    Random Forest (Embedded): Prune low-importance features and optionally train an FCN environment model.
  3. 3.
    Backward Stepwise (Wrapper): Further refine the most energy-critical features using OLS-based feature elimination.
  4. 4.
    Filter-Based (Correlation): Analyze temperature correlations (per-core or per-cluster) to guide thermal-aware scheduling.
  5. 5.
    Task Allocation: Prioritize assigning tasks to the least-correlated units, mitigating overheating and reducing energy consumption.
By unifying these steps, we handle both fine-grained and coarse-grained thermal sensor inputs without losing accuracy or tractability. Random Forest algorithm provides the dimensionality reduction required for backward stepwise selection to reduce its computational overhead and their combination ensures only the most relevant features to the energy consumption remain. The filter-based approach on temperature readings further addresses the correlation of the cores to each other for core allocation to minimize the thermal hotspots, offering a novel, scalable solution validated across platforms (e.g., 61.6% MSE improvement over SOTA).
4 Experiments
This section presents the experimental setup, implementation, and results of our generalized task-to-core allocation framework, integrating Random Forest (RF) feature reduction, backward stepwise OLS selection, and filter-based temperature correlation. To ensure novelty beyond prior filter-based approaches, we validate this hybrid methodology across diverse platforms, demonstrating its effectiveness in optimizing energy and thermal behavior. For clarity and broad applicability, we detail platforms, benchmarks, and metrics, extending beyond Intel-specific results to cluster-based systems. Rigorous quantitative outcomes—backed by multi-platform experiments and comparisons to state-of-the-art (SOTA)—establish the framework’s soundness. We outline the setup, training, and empirical findings below, linking results to the methodology’s advanced feature selection and modeling strategies.
4.1 Experimental Platform, Benchmarks, and Evaluation
All results and figures primarily reflect experiments on a superscalar Intel Core i7 12th Gen with 14 cores (8 P-cores, 6 E-cores, per-core temperature via sensors), chosen for its hybrid architecture. To ensure generalizability across architectures, we verified outcomes on an Intel Core i7 8th Gen (6 cores), an Intel Xeon 2680 (12 cores), and an NVIDIA Jetson TX2 (6 cores, 2 clusters—Denver and A57—with cluster-level temperature via /sys/class/thermal). This multi-platform approach validates the framework’s adaptability to per-core and cluster-based systems, addressing the need for broader applicability.
We evaluated each platform using the Barcelona OpenMP Tasks (BOTS) suite Duran et al. (2009), featuring diverse parallel workloads (e.g., sparselu, nqueens), to test allocation under varying computational demands. Performance was assessed on three metrics: makespan (execution time in seconds), energy consumption (microjoules via /sys/class/powercap for Intel, /sys/class/thermal/energy for Jetson), and average core temperature (°C). These metrics capture performance, power, and thermal trade-offs, ensuring practical impact. Each BOTS benchmark ran 10 times per configuration, with results averaged to reduce variability. A temperature cooldown threshold of 70°C between runs ensured thermal stability, enhancing experimental rigor.
4.2 Implementation and Training Details
The framework was implemented in Python, leveraging subprocess for system calls (e.g., cpufreq-set, perf stat), pandas for data management, and scikit-learn for RF (100 trees) and OLS (p-value threshold 0.05). Parsl facilitated parallel task execution, scaling across platforms. Data from each platform’s profiler and temperature sensors were split into 80% training and 20% test sets. Key hyperparameters—batch size (32), hidden neurons (64–256), epochs (50–200), and learning rate (0.001–0.01)—were tuned via grid search, balancing convergence and generalization. Mean Squared Error (MSE) served as the primary loss criterion, ensuring predictive accuracy.
We evaluated multiple neural network architectures—Fully Connected Networks (FCN), Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM) networks, Convolutional Neural Networks (CNN), and Attention-based models (e.g., Transformers)—retaining top performers based on validation loss. Feature subsets were derived using the methodology’s three stages: Pearson correlation for temperature dependencies, backward stepwise OLS for energy-correlated features, and RF importance rankings for reduction. Bootstrapping (100 resamples) reduced overfitting and increased robustness, preserving critical predictors while providing variance insights under different sampling scenarios, contributing to the framework’s statistical rigor.
4.3 Empirical Results
We developed two predictive models: a profiler model for energy consumption and performance metrics (e.g., cache miss rates, branch miss rates, CPU cycles, instructions per cycle, average speed, page faults, context switches) and a temperature model for future thermal behavior. The profiler model incorporated all available features (e.g., 75 on Intel Core i7 12th Gen) plus current per-core temperatures $\theta_{i}$ and differences $\Delta\theta_{i}=\theta_{i,t}-\theta_{i,t-1}$. The temperature model used 30 features on Intel Core i7 12th Gen (14 temperatures, 14 differences, average temperature), scaled to 6 features on Jetson TX2 (2 cluster temperatures, 2 differences, average), maintaining diversity without overhead.
Figures 4, 6, 3(a), and 3(b) showcase Intel Core i7 12th Gen results, demonstrating that fewer than 8 features (e.g., frequency, cache misses) suffice for accurate predictions, reducing computational costs. Figure 2 shows RF-based feature selection using 39 out of 72 predictors for profiler data and 21 out of 31 for temperature yields nearly optimal CV error, highlighting efficiency critical for real-time embedded systems with strict power and thermal budgets.
The algorithm dynamically adapts to thermal behavior, distributing load evenly across cores or clusters. Figure 6 compares correlation-based (Corr) and random (Rand) core selection under different governors on Intel Core i7 12th Gen. While random allocation leverages unbiased distribution, correlation-based allocation excels in subset scenarios (e.g., 10 of 14 cores). Multi-platform validation reinforces these trends across Jetson TX2, Intel Xeon 2680, and Intel Core i7 8th Gen, proving the approach’s adaptability beyond a single architecture.
Figure 5: Comparison of average temperature and energy consumption for correlation-based (Corr) and random (Rand) core selection. Experiments performed on Intel Core i7 12th Gen processor with 14 cores.
Figure 6: Comparison of FCN models with and without feature selection and bootstrapping on Intel Core i7 12th Gen.
During training, hyperparameters were tuned to optimize convergence speed and generalization, with models saved at optimal checkpoints (e.g., when validation error plateaued or declined sharply). Independent neural networks for temperature and profiler tasks ensured specialized performance, delivering precise predictions tailored to each objective.
Table 1 lists top features from Random Forest (RF) and backward stepwise OLS selections for Intel Core i7 12th Gen. Tables 2 and 3 present regression statistics for temperature and profiler predictors against energy consumption.
Dataset Random Forest (Top Features) Stepwise OLS (Top Features)
[…]
Table 1 shows RF selecting a broader feature set (e.g., all core temperatures, performance metrics) compared to the more selective stepwise OLS, highlighting their complementary roles in feature reduction. For the profiler dataset, Task ID consistently ranks high in both methods, underscoring its key role in identifying energy consumption patterns, followed by Temp Core 1 and Temp Core 13 in OLS as significant predictors. This suggests that cores 1 and 13, when heavily utilized, may disproportionately influence energy use due to their temperature profiles. Consequently, an allocation strategy minimizing task assignments to these cores could enhance energy efficiency, a hypothesis supported by their prominence in the feature selection process.
Xeon (12 cores) Core i7 8th Gen (4 cores) Core i7 12th Gen (14 cores)
$\theta_{i}$ Est. t-val. p-val. $\theta_{i}$ Est. t-val. p-val. $\theta_{i}$ Est. t-val. p-val.
$\theta_{1}$ 2.35e5 5.76 8.48e-9 $\theta_{0}$ 3.64e3 5.69 1.31e-8 $\theta_{2}$ -3.79e3 -6.83 8.41e-12
$\theta_{9}$ -2.53e5 -5.38 7.38e-8 $\theta_{1}$ -9.77e3 -4.18 2.87e-5 $\theta_{0}$ 4.87e2 2.76 5.81e-3
$\theta_{7}$ -1.56e5 -3.10 1.93e-3 $\theta_{3}$ 8.00e3 3.57 3.64e-4 $\theta_{3}$ -5.96e3 -2.05 4.06e-2
$\theta_{6}$ -9.90e4 -2.04 4.09e-2 $\theta_{2}$ -4.63e2 -0.26 7.97e-1 $\theta_{9}$ -2.15e4 -1.93 5.38e-2
$\theta_{8}$ -9.15e4 -1.91 5.66e-2 $\theta_{1}$ 4.77e3 1.88 5.98e-2
$\theta_{4}$ 6.84e4 1.64 1.01e-1 $\theta_{13}$ -1.22e4 -1.77 7.62e-2
$\theta_{2}$ 6.62e4 1.59 1.12e-1 $\theta_{10}$ 9.45e3 1.47 1.41e-1
$\theta_{11}$ -4.82e4 -1.10 2.72e-1 $\theta_{4}$ 3.43e3 1.02 3.10e-1
$\theta_{10}$ -2.29e4 -0.52 6.02e-1 $\theta_{11}$ 6.64e3 0.91 3.65e-1
$\theta_{0}$ -2.04e4 -0.52 6.02e-1 $\theta_{7}$ 6.87e3 0.43 6.67e-1
$\theta_{3}$ -4.50e3 -0.10 9.18e-1 $\theta_{5}$ 1.40e3 0.38 7.05e-1
$\theta_{5}$ 4.68e1 0.00 9.99e-1 $\theta_{12}$ 1.06e3 0.13 8.94e-1
$\theta_{6}$ 8.30e2 0.07 9.44e-1
$\theta_{8}$ 8.11e1 0.01 9.96e-1
Table 2: Temperature predictors ($\theta_{i}$) vs. energy: Estimates (Est.), t-values (t-val.), p-values (p-val.), sorted by p-value. Significant predictors (e.g., $\theta_{1}$, $\theta_{0}$, $\theta_{2}$) show strong energy impact.
Table 2 evaluates per-core temperatures’ effect on energy, with low p-values indicating significant predictors across platforms (e.g., $\theta_{1}$ on Xeon, $\theta_{2}$ on Core i7 12th Gen).
Xeon (12 cores) Core i7 8th Gen (4 cores) Core i7 12th Gen (14 cores)
Predictor Est. t-val. p-val. Predictor Est. t-val. p-val. Predictor Est. t-val. p-val.
Context Switch Rate -3.07e4 -34.90 1.34e-261 Elapsed Time 4.85e7 39.51 0.00e0 Elapsed Time 4.67e7 35.82 3.29e-275
Cache Miss % Refs 3.32e6 34.45 4.02e-255 Cache Miss 2.23e5 21.35 3.80e-100 Atom Branch Misses 1.41e0 25.10 1.21e-137
Cache Refs 2.82e-1 33.53 4.94e-242 Cache Refs 6.39e-2 16.03 1.48e-57 Cache Misses 4.18e-1 24.65 7.37e-133
Elapsed Time 6.44e7 31.52 1.52e-214 Total Cores -5.34e6 -15.47 1.09e-53 Branch Misses 1.36e0 22.78 6.66e-114
Cycles -1.98e-1 -16.52 5.22e-61 Branch Miss 3.37e6 10.41 2.43e-25 Atom Branches -1.93e-2 -18.87 5.72e-79
Branches -1.69e-1 -14.19 1.59e-45 Instructions -1.48e-3 -6.51 7.61e-11 Atom Cycles Freq -2.31e6 -17.85 7.14e-71
CPU Cycles -4.33e-2 -10.94 8.14e-28 Context Switch Rate -2.43e2 -5.42 6.10e-8 Atom Branches Rate 7.26e3 17.30 1.06e-66
Context Switches -2.14e3 -10.15 3.53e-24 Context Switches 2.07e3 5.22 1.82e-7 Branches 1.17e4 15.79 6.30e-56
Cache Misses -4.37e-1 -9.78 1.47e-22 Insn per Cycle 5.84e5 4.33 1.52e-5 Atom Cycles -1.07e-1 -15.33 7.10e-53
Total Cores 7.82e4 9.69 3.43e-22 Branch Misses -1.06e0 -4.18 2.95e-5 User Time -1.31e8 -11.71 1.33e-31
Cycles Freq 1.74e8 9.65 5.41e-22 Cache Misses 1.04e-1 3.88 1.03e-4 Sys Time -1.29e8 -11.59 5.34e-31
CPU Clock ms 2.63e6 8.81 1.29e-18 CPU Clock -5.82e5 -3.68 2.30e-4 Faults 3.77e2 11.31 1.32e-29
User Time -2.27e8 -8.63 6.37e-18 CPU Cycles Freq 1.69e5 3.56 3.65e-4 Cycles Freq 7.57e4 11.31 1.40e-29
Sys Time -2.20e8 -8.38 5.60e-17 Task Clock 5.60e5 3.55 3.91e-4 Page Faults -3.70e2 -11.23 3.42e-29
Cache Refs Rate 6.97e5 8.05 8.51e-16 CPU Cycles 6.89e-4 2.64 8.18e-3 Avg Freq 1.14e3 10.39 3.00e-25
Task Clock -1.80e6 -6.17 7.12e-10 Page Fault Rate 9.68e3 2.45 1.41e-2 Atom Cache Miss -1.29e4 -10.15 3.74e-24
Branch Misses 1.82e0 4.24 2.22e-5 Fault Rate 9.68e3 2.45 1.42e-2 Atom Branch Miss 9.16e0 9.09 1.07e-19
Insn per Cycle -9.52e6 -4.23 2.36e-5 Task Util 9.01e6 2.28 2.25e-2 Instructions -4.71e-4 -8.74 2.45e-18
Task Util -1.41e8 -2.30 2.12e-2 Cycles Freq 3.52e5 1.39 1.64e-1 Cache Miss 1.25e4 7.55 4.56e-14
CPUs Used 1.25e8 2.04 4.14e-2 Branches Rate 1.56e3 1.34 1.80e-1 Branches Rate -1.77e-1 -6.68 2.44e-11
CPU Cycles Freq 2.00e6 1.20 2.30e-1 User Time 4.06e6 1.04 2.99e-1 Cycles -1.11e-3 -5.93 3.04e-9
Page Faults -2.64e1 -1.18 2.36e-1 Cycles 1.15e-3 0.99 3.21e-1 Branch Miss 6.12e0 5.86 4.69e-9
Faults -2.62e1 -1.17 2.41e-1 CPUs Used -3.43e6 -0.87 3.85e-1 CPU Cycles -8.56e-4 -5.71 1.14e-8
Instructions 7.52e-4 0.51 6.09e-1 Sys Time 3.21e6 0.82 4.12e-1 Atom Cycles 1.12e-3 5.67 1.47e-8
Fault Rate -1.69e2 -0.00 9.97e-1 Faults 5.90e0 0.53 5.95e-1 Branch Miss 2.58e4 5.37 8.12e-8
Page Fault Rate -1.69e2 -0.00 9.97e-1 Branches -8.77e-4 -0.17 8.69e-1 Context Switch Rate -5.68e2 -4.86 1.20e-6
Page Faults 1.90e-1 0.02 9.86e-1 Atom Cache Miss 1.09e0 4.61 4.01e-6
Table 3: Profiler predictors (excl. temp, freq, speed) vs. energy: Estimates (Est.), t-values (t-val.), p-values (p-val.), sorted by p-value. Key predictors (e.g., Context Switch Rate, Elapsed Time) drive energy variance.
Table 3 highlights profiler features’ statistical significance, with metrics like Context Switch Rate and Cache Misses showing strong energy correlations across platforms.
4.4 Comparative Model Analysis
We tested multiple neural architectures to identify optimal designs for energy and temperature prediction, benchmarking against a SOTA approach Hosseinimotlagh et al. (2021) to establish novelty and rigor. Table 4 details MSE percentage ($MSE\times 100$) and parameter counts on Intel Core i7 12th Gen. Baseline FCN achieved solid performance (MSE 1.0299 temperature, 3.9047 profiler), but FCN with RF feature selection (FCN+RF) improved accuracy (0.9808, 2.4862) with fewer parameters (1694 vs. 2014 temperature; 2699 vs. 3787 profiler). Adding bootstrapping (FCN+RF+BS) yielded the best results (0.9640, 2.4669), maintaining lightweight models ideal for resource-constrained systems.
Table 4: MSE percentage and total number of parameters for different architectures on Intel Core i7 12th Gen.
Model Temperature Profiler
MSE Params MSE Params
FCN 1.0299 2014 3.9047 3787
FCN+RF 0.9808 1694 2.4862 2699
FCN+RF+BS 0.9640 1694 2.4669 2699
RNN 1.0119 3070 2.8493 4843
LSTM 1.0307 9310 2.7778 15115
Conv 1.0134 5118 2.8217 6891
Attention 1.0143 6238 2.8933 8011
SOTA Hosseinimotlagh et al. (2021) 2.5000 - - -
Figures 8 and 8 present prediction examples from Intel Core i7 12th Gen thermal and profiler data. The FCN model uses the complete feature set, while FCN+RF leverages a reduced subset identified by RF, forecasting environmental behavior with notable similarity to full-feature predictions. This validates RF’s effectiveness in compressing features without sacrificing accuracy, enhancing computational efficiency.
Figure 7: Results for temperature prediction from ground truth on Intel Core i7 12th Gen for regular FCN and FCN via Random Forest feature reduction strategy.
Figure 8: Results for profiler data prediction from ground truth on Intel Core i7 12th Gen for regular FCN and FCN via Random Forest feature reduction strategy.
Figure 6 illustrate the effect of feature selection and bootstrapping on test MSE. Omitting either increases MSE by 20–30%, underscoring their combined benefit. By focusing on a small, influential feature subset, the final FCN models improve prediction accuracy and remain computationally light, aligning with real-time and power constraints in multi-core embedded systems. This synergy of reduced feature sets, resampling, and specialized architectures (e.g., FCN+RF+BS) proves effective across the BOTS workloads and platforms tested, surpassing simpler filter-based methods and establishing the framework’s novelty and practical value.
Overall, the advanced feature selection (filter, wrapper, and embedded) combined with neural network models enhances energy and temperature prediction. Multi-platform results and comparisons to SOTA validate the approach’s effectiveness, rigor, and applicability to diverse multi-core systems.
5 Related Work
Probabilistic Methods for DVFS and Task-to-Core Allocation
Probabilistic worst-case execution time (pWCET) and worst-case energy consumption (pWCEC) estimation in embedded real-time systems leverage measurement-based and static methods Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017). Pallister et al. Pallister et al. (2017) studied instruction-specific impacts on pWCEC, and extreme value theorem (EVT) Edgar and Burns (2001) provided upper bounds for performance metrics. However, these approaches rarely quantify the statistical significance of system metrics for energy or latency bounds. Our work introduces a hybrid statistical learning framework, systematically prioritizing feature importance to optimize energy-efficient DVFS and task-to-core allocation across diverse platforms, surpassing traditional probabilistic bounds.
Statistical Learning
Statistical learning has been employed to pinpoint features for energy optimization and scheduling, often using hardware events or application traits Sasaki et al. (2007); Cazorla et al. (2019); Liu et al. (2021). Sasaki et al. Sasaki et al. (2007) applied decision trees to streamline DVFS table lookups, while Cazorla et al. Cazorla et al. (2019) used hardware counters for energy reduction. Liu et al. Liu et al. (2021) correlated compiler features with latency. Yet, these efforts underexplored runtime metrics and sampling strategies critical for accuracy. We enhance this by integrating runtime performance metrics with a novel feature selection pipeline, validated across Intel Xeon and Core i7 platforms, improving parallel scheduling and environment modeling precision.
Low-Energy DVFS and Task-to-Core Allocation
Low-energy multicore scheduling via DVFS and task-to-core mapping is well-studied Xie et al. (2021, 2017); Zhu et al. (2004); Jiang et al. (2019); Chen et al. (2018); Zhou et al. (2018); Kim and Wu (2020); Dinakarrao et al. (2019); Shen et al. (2012); Wang et al. (2017). Xie et al. Xie et al. (2021) surveyed heuristic and machine learning methods for energy-constrained scheduling. Many such ML approaches, however, require extensive datasets and computational resources, limiting practicality. Our method mitigates these issues with a lightweight statistical model, reducing data needs while maintaining accuracy, as demonstrated on embedded and heterogeneous systems. Hierarchical multi-agent approaches Pivezhandi et al. (2026b) and graph-driven performance models Pivezhandi et al. (2026a) have also shown promise for DAG workloads, offering scalable alternatives to existing high-overhead techniques.
Few-Shot RL
Few-shot learning, including transfer learning and meta-learning, minimizes data demands in scheduling Wang et al. (2020); Lee et al. (2020); Wang et al. (2016); Florensa et al. (2017); Arora and Doshi (2021). MAML Finn et al. (2017) adapts to new tasks with few samples, and model-based RL Moerland et al. (2023) approximates dynamics to limit real-world data use. Works like Lin et al. (2023); Kim et al. (2021); Zhou and Lin (2021); Zhang et al. (2024) underutilize statistical resampling and feature ranking for energy-efficient scheduling. Recent advances include flow-augmented data generation for few-shot RL Pivezhandi and Saifullah (2026a) and zero-shot LLM-guided resource allocation Pivezhandi and Saifullah (2026b). Our approach fills this gap, combining resampling with a robust feature selection strategy, validated across platforms, to enhance few-shot task-to-core allocation efficiency.
Feature Evaluation
Thermal-aware scheduling models often predict behavior using utilization or temperature data to prevent throttling Yan et al. (2003); Brooks et al. (2007); Maity et al. (2022); Hosseinimotlagh and Kim (2019); Singla et al. (2015); Li and Wu (2012); Kassab et al. (2021). Studies like Maity et al. (2022); Lin et al. (2023); Kim et al. (2021) model transients and ambient conditions Hosseinimotlagh et al. (2021), but lack rigorous statistical analysis of feature correlations. We address this by applying a systematic correlation-based feature evaluation, rigorously tested on real hardware, to uncover interdependencies and optimize DVFS and allocation, improving energy and performance outcomes over prior heuristic-driven methods.
6 Conclusion
We demonstrated the effectiveness of feature selection using statistical learning for environment modeling and task-to-core allocation in embedded systems. Our correlation-aware task-to-core allocation reduces energy consumption by up to 10% and temperature by up to 5C compared to random core selection. The compressed bootstrapped regression model reduces thermal prediction error by 6% and the number of parameters by 16%. Tested on Intel Core i7 8th and 12th generation, Intel Xeon 2680 processors and Jetson TX2, our method shows a 61.6% reduction in mean squared error compared to state-of-the-art approach. This finding paves the way for future use of statistical learning methods in performance efficiency of task-to-core allocation in heterogeneous processors.
ACKNOWLEDGEMENT
The work was supported by the US National Science Foundation through grants CNS-2301757, CAREER- 2306486, CNS-2306745, and by the US Office of Naval Research through grant N00014-23-1-2151.